ganeshie8 (ganeshie8):

[Solved by @Kainui @thomas5267 @mathmath333 @ikram002p ] Find one prime divosor of \[\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} + 23\]

OpenStudy (anonymous):

Im going to go ahead and guess 23

OpenStudy (ikram002p):

Did you find any ?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

(i dont want to spoil the fun by saying anything about finding the factor :P)

OpenStudy (kainui):

I think I found one ahhahaha but I have no way of checking.

OpenStudy (kainui):

No wait, yes I do. Yeah I found a factor. I won't say it cause it's funny haha.

ganeshie8 (ganeshie8):

yes kainui got it xD

OpenStudy (anonymous):

probably is the number itself, I bet it's a prime -.^

OpenStudy (ikram002p):

i would assume there is a factor x then \(\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} + 23 \equiv 0 \mod x\) \(\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} \equiv -23 \mod x\)

OpenStudy (thomas5267):

Just trial and error should work.

OpenStudy (thomas5267):

The number is small.

ganeshie8 (ganeshie8):

thomas also got it !

ganeshie8 (ganeshie8):

both have msged me the correct divosor and a legitimate method for proving it..

OpenStudy (thomas5267):

@ikram002p You are thinking in the right direction. Your answer inspired me to come up mine's.

OpenStudy (ikram002p):

i have several answers lol but idk away to make sure of it xD

OpenStudy (ikram002p):

and feeling stupid to post lol since already two of you knows :O

OpenStudy (kainui):

I found two factors other than 1 and the number itself lol =P

OpenStudy (anonymous):

ive no idea how to even approach the question....that number looks bigger than everything that came into existence combined.......

OpenStudy (ikram002p):

:O

OpenStudy (thomas5267):

I can't find any others. One is hard enough for me...

ganeshie8 (ganeshie8):

lol i also thought one factor is good enough for me, but kainui seems to be having another factor... i don't understand his solution fully though il need to go through it few more time.. xD

OpenStudy (ikram002p):

well i thought of this xD \(10^n+23 \) for n which is factor of that power hmm which also sounds does not work :O

ganeshie8 (ganeshie8):

@mathmath333 also got it!

OpenStudy (kainui):

I'll give a hint cause it's how I started. Try to write out the number. =P

OpenStudy (ikram002p):

xD

OpenStudy (thomas5267):

Should the stacked exponent read from top to bottom of left to right?

ganeshie8 (ganeshie8):

@ikram002p got it too

ganeshie8 (ganeshie8):

kainui, ikrram and mathmath claim to be having 2 divosors xD

ganeshie8 (ganeshie8):

ikram and mathmath have correct methods which i can understand easily for their 2 divisors !

OpenStudy (thomas5267):

I can't even write the number out. Too many zeros.

ganeshie8 (ganeshie8):

i bet that number doesnt exist in this universe with that many zeros ! ikram found 3 divosors!

OpenStudy (thomas5267):

Wow! I am thoroughly impressed.

OpenStudy (ikram002p):

last one was a joke :P

OpenStudy (kainui):

Haha well I guess I will reveal how I found my answers unless people are still wanting to try to figure it out.

OpenStudy (ikram002p):

ok reveal :O

OpenStudy (ikram002p):

i wanna see @thomas5267 solution though

OpenStudy (thomas5267):

\[ 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} + 23\equiv1^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}}-1\equiv0\pmod3 \]

OpenStudy (ikram002p):

nice, any others ?

OpenStudy (thomas5267):

No I can't think of any other. 11 is not one of the prime divisor though.

OpenStudy (kainui):

Ok first one I found by trying to write it out, I got 100...0023 then I saw that if you add up all the digits you get 6, so it's divisible by 3. Second one I just divided this number by 3. How did I do that? Well just like this number has 5 nines \[\Large 10^5-1=99999\] I noticed that we could rewrite the number as \[\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}}-1 + 24\] So really there are \[10^{10^{10^{10^{10^{10^{10^{10}}}}}}} \text{ nines}\] added to 24, so divide this by 3: \[(999...999 + 24)/3 = 333...333+8\] so Now there are the same number of 3s as there were nines, but if we add 8 we will get the last two digits change to 41 so really there are \[ \left( 10^{10^{10^{10^{10^{10^{10^{10}}}}}}}-2 \right) \text{ threes}\] followed by 41, so you could write it as: \[\Large 333...33341\] so that's the other number. Hahaha. =P

OpenStudy (ikram002p):

:O

ganeshie8 (ganeshie8):

brilliant xD thats another valid factor so all in all we have found four divisors so far xD but only me knows all four factos haha!

OpenStudy (kainui):

Haha is the other one 13 and the number divided by 13?

ganeshie8 (ganeshie8):

thats @mathmath333 's divosor :)

ganeshie8 (ganeshie8):

3 and 13 give us another divosr : 39 xD but yeah 39 is not so primy..

OpenStudy (ikram002p):

i told u that was a joke :P

OpenStudy (kainui):

I wonder if we can try to do some sort of trick to divide the number by 13 too.

OpenStudy (ikram002p):

can we find factors of \[\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} +1\]

ganeshie8 (ganeshie8):

unthinkable

ganeshie8 (ganeshie8):

have some respect for large numbers ikram :P

OpenStudy (kainui):

Yeah seriously what's wrong with you @ikram002p ?

ganeshie8 (ganeshie8):

lol

OpenStudy (kainui):

I know a factor for this number where p is all primes greater than 23 \[\large 10^{10^{10^{10^{10^{10^{10^{10^{10}}}}}}}} +p+1\]

OpenStudy (kainui):

BEAT THAT!

ganeshie8 (ganeshie8):

2 ?

ganeshie8 (ganeshie8):

2| (p+1) 2|10^n so..

OpenStudy (kainui):

Yeah, but I also know 2 more numbers that are factors of that number... **cough**

OpenStudy (kainui):

1, 2, and the number itself. =P

ganeshie8 (ganeshie8):

i thought so.. leme think :)

ganeshie8 (ganeshie8):

lol okay

OpenStudy (ikram002p):

:P

OpenStudy (ikram002p):

4

OpenStudy (ikram002p):

for some p's

OpenStudy (anonymous):

Wow.what the heck?

mathslover (mathslover):

@FrownUpsideDown8767 - Welcome to Mathematics my friend :D

OpenStudy (ikram002p):

ok ok post some other question :D

ganeshie8 (ganeshie8):

im still working on proving 13 is a factor of the number in main question

ganeshie8 (ganeshie8):

so the question is not dead yet ikky :P

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\dfrac{10^{10^n}+23}{13}\hspace{.33em}\\~\\ =\dfrac{((10^6) (10^4))^n+10}{13} \hspace{.33em}\\~\\ =\dfrac{((1)\cdot 100\cdot 100))^n-3}{13}\hspace{.33em}\\~\\ =\dfrac{((1)\cdot (81))^n-3}{13}\hspace{.33em}\\~\\ =\dfrac{3^{n~~ mod ~~3}-3}{13}\\ \normalsize {as~~ 3^3 mod ~~13=1} \hspace{.33em}\\~\\ =\dfrac{3^1-3}{13} \hspace{.33em}\\~\\ =0 }\end{align}\)

OpenStudy (ikram002p):

hahahaha @mathmath333 did it

OpenStudy (thomas5267):

mathmath333 answer is so difficult to understand...

OpenStudy (mathmath333):

note that i used the cycle of mod 13 againts 3

OpenStudy (thomas5267):

I can't even understand your second step. Could you explain it?

OpenStudy (mathmath333):

the cycle of 3^\(\bf3\)=27 =1 mod 13) so it is showing the cycle of 3 that why i applied mod 3 on \(n\)

OpenStudy (mathmath333):

this works in the cases for \(\huge \begin{align} \color{black}{a^{b^c}\ }\end{align}\)

OpenStudy (mathmath333):

this stuff we do to find the remainders

ganeshie8 (ganeshie8):

mathmath's proof hinges on thae fact that we can write \[\large 10^{10^{\cdots}} = 10^{10k} = (10^{10})^k\] for some integer \(k\)

OpenStudy (thomas5267):

How does \( \dfrac{10^{10^n}+23}{13}=\dfrac{((10^6) (10^4))^n+10}{13} \)?

ganeshie8 (ganeshie8):

thats not true, below is true how ever \[\dfrac{10^{10n}+23}{13}=\dfrac{((10^6) (10^4))^n+10}{13}\]

ganeshie8 (ganeshie8):

i dont see any flaw in rest of his work..

OpenStudy (mathmath333):

yep thats true

OpenStudy (mathmate):

and the other factor is 769230769230......769230771 since 13|10023, 10000000023, 1000000000000023.... with quotients 771, 769230771, 769230769230771, ...

OpenStudy (mathmath333):

i will find if i can given a more elegant proof for this

ganeshie8 (ganeshie8):

thats a very interesting pattern @mathmate XD

OpenStudy (mathmate):

10023/13=771 That's because 13|999999 and the quotient is 76923 and 10000000023-10023=9999990000+10023

ganeshie8 (ganeshie8):

Nice :) period of 1/13 is 6 so 13 | (10^6 - 1)

OpenStudy (mathmate):

\(\dfrac{10^(6k+2)+23}{13}=\dfrac{10(6k)-1)*10000+10023}{13}=0~ mod~ 13\)

OpenStudy (ikram002p):

i have other solution though but seem way as mathmath thought off

OpenStudy (mathmate):

Finally get it right: \(\dfrac{10^{6k+2}+23}{13}=\dfrac{(10^{6k}-1)*10000+10023}{13}=0~ mod~ 13\)

OpenStudy (ikram002p):

10=-3 mod 13 (-3)^12=1 mod 13 (-3)^12n=1 mod 13 (-3)^12n+1=-3 mod 13 (-3)^12n+2=9 mod 13 (-3)^12n+3=-1 mod 13 (-3)^12n+4=3 mod 13 ........ so we need to show 10^stuff is of the form 12n+4

OpenStudy (anonymous):

@mathmath333 is a boss!!!

OpenStudy (thomas5267):

What happened here? \[ \dfrac{((1)\cdot (81))^n-3}{13} =\dfrac{3^{n\text{ mod }3}-3}{13} \]

OpenStudy (thomas5267):

God I understand now. Definitely one of the hardest proof I have ever read.

OpenStudy (mathmath333):

i m trying to come up with another proof

OpenStudy (thomas5267):

I will try to make his proof more understandable.